# Qumode Tensor Networks for False Vacuum Decay in Quantum Field Theory

### v1

@@ II.3 Non-renormalisable theories: the double $\tanh(\varphi^{\ell})$ well potential
${\mathscr{V}}(\varphi)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \frac{\lambda}{\ell!}\tanh(\mu^{-\ell}(\varphi-\varphi_{\rm fv})^{\ell})$ (24)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\leavevmode\nobreak\ \left(\frac{\lambda}{\ell!}+\varepsilon\right)\left(\tanh(\mu^{-\ell}(\varphi-\varphi_{\rm tv})^{\ell})-1\right)\leavevmode\nobreak\ ,$
@@
${\mathscr{V}}(\varphi)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \frac{\lambda}{\ell!}(\varphi-\varphi_{\rm fv})^{\ell}-\frac{\lambda}{3\ell!}(\varphi-\varphi_{\rm fv})^{3\ell}\leavevmode\nobreak\ +\leavevmode\nobreak\ \ldots$ (25)
@@ II.4 Solvable non-renormalizable qumodes: the Pöschl-Teller potential
${\mathscr{V}}(\varphi)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \frac{\alpha^{2}\gamma(\gamma+1)}{2}\tanh^{2}(\alpha(\varphi-\varphi_{\rm fv}))$ (26)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\leavevmode\nobreak\ \left(\frac{\alpha^{2}\gamma(\gamma+1)}{2}+\varepsilon\right){\rm sech}^{2}(\alpha(\varphi-\varphi_{\rm tv}))\leavevmode\nobreak\ .$
@@
${\mathscr{V}}_{\rm PT}(\varphi)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \frac{\alpha^{2}\gamma(\gamma+1)}{2}\tanh^{2}(\alpha(\varphi-\varphi_{\rm fv}))\leavevmode\nobreak\ .$ (27)
@@
$\psi_{n}^{(0)}(q_{n})\leavevmode\nobreak\ =\leavevmode\nobreak\ \langle q_{n}|0\rangle_{-}\leavevmode\nobreak\ =\leavevmode\nobreak\ {\cal N}_{0}\leavevmode\nobreak\ {\rm sech}^{\gamma}(\alpha(q_{n}-\varphi_{\rm fv}))$ (28)
@@
$\mathscr{E}^{-}_{0}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\alpha^{2}}{2}\gamma\leavevmode\nobreak\ ,$ (29)
@@
${\cal N}_{0}\leavevmode\nobreak\ =\leavevmode\nobreak\ {\sqrt{\frac{\alpha}{\sqrt{\pi}}\frac{\,\Gamma\left(\gamma+\frac{1}{2}\right)}{\leavevmode\nobreak\ \Gamma(\gamma)}}}\leavevmode\nobreak\ .$ (30)
@@
$\psi_{n}^{(\nu)}(q_{n})\leavevmode\nobreak\ =\leavevmode\nobreak\ \langle q_{n}|\nu\rangle\leavevmode\nobreak\ =\leavevmode\nobreak\ {\cal N}^{\gamma-\nu}_{\gamma}P^{\gamma-\nu}_{\gamma}(\tanh(\alpha(q_{n}-\varphi_{\rm fv})))\leavevmode\nobreak\ ,$ (31)
@@
$\mathscr{E}^{-}_{\nu}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\alpha^{2}}{2}\left(\gamma(\gamma+1)-(\gamma-\nu)^{2}\right)\leavevmode\nobreak\ .$ (32)
@@
$\mathscr{E}^{+}_{\nu}\leavevmode\nobreak$ $=\leavevmode\nobreak\ \frac{\alpha^{2}}{2}\left((\gamma+1)(\gamma+2)-(\gamma+1-\nu)^{2}\right)\leavevmode\nobreak\ -\varepsilon\leavevmode\nobreak$
$=\leavevmode\nobreak\ \frac{\alpha^{2}}{2}\left(\gamma(\gamma+1)-(\gamma+1-\nu)^{2}\right)\leavevmode\nobreak\ \leavevmode\nobreak\ .$ (33)
@@
$|{\mathscr{E}}_{1}\rangle\leavevmode\nobreak$ $=\leavevmode\nobreak\ \cos\theta|0\rangle_{-}+\sin\theta|1\rangle_{+}$
$|{\mathscr{E}}^{\prime}_{1}\rangle\leavevmode\nobreak$ $=\leavevmode\nobreak\ -\sin\theta|0\rangle_{-}+\cos\theta|1\rangle_{+}\leavevmode\nobreak\ ,$ (34)
@@
$|\psi_{n}(t)\rangle\leavevmode\nobreak$ $=\leavevmode\nobreak\ \cos(\Delta\mathscr{E}\,t/2)\,|0\rangle_{-}$ (35)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\leavevmode\nobreak\ i\sin(\Delta\mathscr{E}\,t/2)\,\left(\cos 2\theta|0\rangle_{-}+\sin 2\theta|1\rangle_{+}\right)\leavevmode\nobreak$
@@
$\Delta\mathscr{E}\leavevmode\nobreak\ \sim\leavevmode\nobreak\ {\mathscr{E}}_{1}e^{-S_{1}}$ (36)
@@
[…]
@@ II.5 The qumode lattice
${\mathcal{U}}_{n}(\delta t)\leavevmode\nobreak\ =\leavevmode\nobreak\ e^{-i\hat{p}_{n}^{2}\delta t}e^{-i\left(\hat{q}_{n}^{2}+\mathscr{V}(\hat{q}_{n})\right)\delta t}\leavevmode\nobreak\ ,$ (38)
@@
${\mathcal{U}}_{\rm diag}(\delta t)\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{n=1}^{N}{\mathcal{U}}_{n}(\delta t)\leavevmode\nobreak\ .$ (39)
@@
$\mathcal{U}_{\textrm{hop}}(\delta t)\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{n=1}^{N}e^{i\,\hat{q}\,_{{n+1}}\,\hat{q}_{n}\,\delta t}\leavevmode\nobreak$ (40)
@@
${\cal U}(\delta t)\leavevmode\nobreak\ =\leavevmode\nobreak\ \mathcal{U}_{\textrm{diag}}(\delta t)\leavevmode\nobreak\ \mathcal{U}_{\textrm{hop}}(\delta t)$ (41)
@@
${\cal U}_{\rm exact}(t)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \left[{\cal U}(\delta t)\right]^{t/\delta t}+{\cal O}(\delta t^{2})$ (42)
@@
${\cal U}(\delta t)\leavevmode\nobreak\ =\leavevmode\nobreak\ \mathcal{U}_{\textrm{diag}}(\delta t/2)\leavevmode\nobreak\ \mathcal{U}_{\textrm{hop}}(\delta t/2)\leavevmode\nobreak\ \mathcal{U}^{T}_{\textrm{hop}}(\delta t/2)\leavevmode\nobreak\ \mathcal{U}_{\textrm{diag}}(\delta t/2)\leavevmode\nobreak\ ,$ (43)
@@
${\cal U}_{\rm exact}(t)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \left[{\cal U}(\delta t)\right]^{t/\delta t}+{\cal O}(\delta t^{3})\leavevmode\nobreak\ .$ (44)
@@
II.6 Initial state preparation
Finally, we turn to the initial state preparation for the entire qumode lattice. We have seen that it is possible to determine the “false vacuum ground state” of a single uncoupled anharmonic qumode using the adiabatic theorem. However the ground state of the QFT is not simply the product of ground states of the uncoupled anharmonic qumode oscillators. As discussed in Ref. [1] the QFT ground state approaches the product of uncoupled qumode ground states for modes in the large mass and low momentum limit, but generally it will involve multi-qumode entanglement and hence long-range correlations. This distinction could be ignored in Ref. [1] because that work was concerned only with scattering processes, so it could accordingly focus on scattering of low momentum wave-packets. Unfortunately for tunnelling processes the distinction is important, because the long-range fluctuations that seed the process are absent from the product of qumode groundstates. In other words, quantum tunnelling could occur on the lattice spontaneously even if it is initialised as the product of qumode ground states, but it could only be a much rarer essentially zero momentum coherent transition rather than a localised bubble.
In principle one can initialise the entire lattice using the same adiabatic evolution procedure we used for a single qumode in Sec. II.2.
In detail, we begin the system in an MPS which consists of the product of qumodes all in their “false vacuum ground states” determined as described in II.2. Then we act on the MPS with the following set of $M$ MPOs:
$\widetilde{\mathcal{W}}\leavevmode\nobreak\ =\leavevmode\nobreak\ \bigotimes_{r=1}^{M}\prod_{n=1}^{N}\left[e^{-i\hat{p}_{n}^{2}\delta t}e^{-i({\mathscr{V}}(\hat{q}_{n})+{\mathscr{V}}_{\rm lift}(\hat{q}_{n}))\delta t}\leavevmode\nobreak\ {\widetilde{\mathcal{V}}}^{\,r}_{n}\right]\leavevmode\nobreak\ ,$ (47)
@@ II.6 Initial state preparation
$\widetilde{\mathcal{V}}_{n<N}\leavevmode\nobreak\ =\leavevmode\nobreak\ e^{-i(\hat{q}_{n}^{2}-\hat{q}_{n}\hat{q}_{n+1})\delta s\,\delta t}\leavevmode\nobreak\ ,$ (48)
@@
The principle is, of course, exactly the same as for the determination of the “false vacuum ground state” of a single qumode. That is, beginning in an MPS which is a product of qumode oscillator ground states, and adiabatically turning on the space-derivative couplings, which include the hopping terms, should result in the system reaching the “false vacuum ground state” of the QFT. Note, for this purpose that we have added ${\mathscr{V}}_{\rm lift}$ to the potential to ensure that the system does not tunnel during the adiabatic evolution while it is being prepared. Note also that one could combine everything into a single adiabatic procedure, starting in an MPS which is a product of SHO ground states centred at $\varphi_{\rm fv}$, and adiabatically turning on both ${\mathscr{V}}_{1}$ and the space-derivative terms at the same time.
This kind of procedure would certainly be the initialisation of choice for a qumode lattice built on a genuine continuous-variable quantum computer. However, on a tensor network this procedure would be quite lengthy to implement, and so we will leave it as an exercise for future work.
@@
The alternative method that we will make use of here is to instead approximate the QFT ground state by sampling. That is, one can generate an ensemble (labelled by $s$) of initial position and momentum fluctuations, $\{\Delta q_{n}^{(s)},\Delta p_{n}^{(s)}\}$, whose distributions incorporate those quantum fluctuations and correlations of the QFT ground state which are missing when we simply start the system as an MPS which is a product of quantum anharmonic oscillators. The derivation of these fluctuations is given in App. B. The upshot is that we can mimick the required long-range correlations by setting the qumodes in their “false vacuum ground states” and then augmenting them by random sampling the classical distributions of position and momentum fluctuations presented in Eq. (96), which we reproduce here for convenience:

### v2

@@ II.3 Non-renormalisable theories: the double $\tanh(\varphi^{\ell})$ well potential
${\mathscr{V}}(\varphi)\penalty 10000\$$=\penalty 10000\ \frac{\lambda}{\ell!}\tanh(\mu^{-\ell}(\varphi-\varphi_{\rm fv})^{\ell})$ (24)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +\penalty 10000\ \left(\frac{\lambda}{\ell!}+\varepsilon\right)\left(\tanh(\mu^{-\ell}(\varphi-\varphi_{\rm tv})^{\ell})-1\right)\penalty 10000\ ,$
@@
${\mathscr{V}}(\varphi)\penalty 10000\$$=\penalty 10000\ \frac{\lambda}{\ell!}(\varphi-\varphi_{\rm fv})^{\ell}-\frac{\lambda}{3\ell!}(\varphi-\varphi_{\rm fv})^{3\ell}\penalty 10000\ +\penalty 10000\ \ldots$ (25)
@@ II.4 Solvable non-renormalizable qumodes: the Pöschl-Teller potential
${\mathscr{V}}(\varphi)\penalty 10000\$$=\penalty 10000\ \frac{\alpha^{2}\gamma(\gamma+1)}{2}\tanh^{2}(\alpha(\varphi-\varphi_{\rm fv}))$ (26)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +\penalty 10000\ \left(\frac{\alpha^{2}\gamma(\gamma+1)}{2}+\varepsilon\right){\rm sech}^{2}(\alpha(\varphi-\varphi_{\rm tv}))\penalty 10000\ .$
@@
${\mathscr{V}}_{\rm PT}(\varphi)\penalty 10000\$$=\penalty 10000\ \frac{\alpha^{2}\gamma(\gamma+1)}{2}\tanh^{2}(\alpha(\varphi-\varphi_{\rm fv}))\penalty 10000\ .$ (27)
@@
$\psi_{n}^{(0)}(q_{n})\penalty 10000\ =\penalty 10000\ \langle q_{n}|0\rangle_{-}\penalty 10000\ =\penalty 10000\ {\cal N}_{0}\penalty 10000\ {\rm sech}^{\gamma}(\alpha(q_{n}-\varphi_{\rm fv}))$ (28)
@@
$\mathscr{E}^{-}_{0}\penalty 10000\ =\penalty 10000\ \frac{\alpha^{2}}{2}\gamma\penalty 10000\ ,$ (29)
@@
${\cal N}_{0}\penalty 10000\ =\penalty 10000\ {\sqrt{\frac{\alpha}{\sqrt{\pi}}\frac{\,\Gamma\left(\gamma+\frac{1}{2}\right)}{\penalty 10000\ \Gamma(\gamma)}}}\penalty 10000\ .$ (30)
@@
$\psi_{n}^{(\nu)}(q_{n})\penalty 10000\ =\penalty 10000\ \langle q_{n}|\nu\rangle\penalty 10000\ =\penalty 10000\ {\cal N}^{\gamma-\nu}_{\gamma}P^{\gamma-\nu}_{\gamma}(\tanh(\alpha(q_{n}-\varphi_{\rm fv})))\penalty 10000\ ,$ (31)
@@
$\mathscr{E}^{-}_{\nu}\penalty 10000\ =\penalty 10000\ \frac{\alpha^{2}}{2}\left(\gamma(\gamma+1)-(\gamma-\nu)^{2}\right)\penalty 10000\ .$ (32)
@@
$\mathscr{E}^{+}_{\nu}\penalty 10000\$$=\penalty 10000\ \frac{\alpha^{2}}{2}\left((\gamma+1)(\gamma+2)-(\gamma+1-\nu)^{2}\right)\penalty 10000\ -\varepsilon\penalty 10000\$
$=\penalty 10000\ \frac{\alpha^{2}}{2}\left(\gamma(\gamma+1)-(\gamma+1-\nu)^{2}\right)\penalty 10000\ \penalty 10000\ .$ (33)
@@
$|{\mathscr{E}}_{1}\rangle\penalty 10000\$$=\penalty 10000\ \cos\theta|0\rangle_{-}+\sin\theta|1\rangle_{+}$
$|{\mathscr{E}}^{\prime}_{1}\rangle\penalty 10000\$$=\penalty 10000\ -\sin\theta|0\rangle_{-}+\cos\theta|1\rangle_{+}\penalty 10000\ ,$ (34)
@@
$|\psi_{n}(t)\rangle\penalty 10000\$$=\penalty 10000\ \cos(\Delta\mathscr{E}\,t/2)\,|0\rangle_{-}$ (35)
$\penalty 10000\ \penalty 10000\ \penalty 10000\ +\penalty 10000\ i\sin(\Delta\mathscr{E}\,t/2)\,\left(\cos 2\theta|0\rangle_{-}+\sin 2\theta|1\rangle_{+}\right)\penalty 10000\$
@@
$\Delta\mathscr{E}\penalty 10000\ \sim\penalty 10000\ {\mathscr{E}}_{1}e^{-S_{1}}$ (36)
@@
$S_{1}\penalty 10000\ \approx\penalty 10000\ \int_{\varphi_{\rm fv}}^{\varphi_{\rm esc}}\sqrt{2{\mathscr{V}(q_{n})}}\penalty 10000\ dq_{n}\penalty 10000\ .$ (37)
@@ II.5 The qumode lattice
${\mathcal{U}}_{n}(\delta t)\penalty 10000\ =\penalty 10000\ e^{-i\hat{p}_{n}^{2}\delta t}e^{-i\left(\hat{q}_{n}^{2}+\mathscr{V}(\hat{q}_{n})\right)\delta t}\penalty 10000\ ,$ (38)
@@
${\mathcal{U}}_{\rm diag}(\delta t)\penalty 10000\ =\penalty 10000\ \prod_{n=1}^{N}{\mathcal{U}}_{n}(\delta t)\penalty 10000\ .$ (39)
@@
$\mathcal{U}_{\textrm{hop}}(\delta t)\penalty 10000\ =\penalty 10000\ \prod_{n=1}^{N}e^{i\,\hat{q}\,_{{n+1}}\,\hat{q}_{n}\,\delta t}\penalty 10000\$(40)
@@
${\cal U}(\delta t)\penalty 10000\ =\penalty 10000\ \mathcal{U}_{\textrm{diag}}(\delta t)\penalty 10000\ \mathcal{U}_{\textrm{hop}}(\delta t)$ (41)
@@
${\cal U}_{\rm exact}(t)\penalty 10000\$$=\penalty 10000\ \left[{\cal U}(\delta t)\right]^{t/\delta t}+{\cal O}(\delta t^{2})$ (42)
@@
${\cal U}(\delta t)\penalty 10000\ =\penalty 10000\ \mathcal{U}_{\textrm{diag}}(\delta t/2)\penalty 10000\ \mathcal{U}_{\textrm{hop}}(\delta t/2)\penalty 10000\ \mathcal{U}^{T}_{\textrm{hop}}(\delta t/2)\penalty 10000\ \mathcal{U}_{\textrm{diag}}(\delta t/2)\penalty 10000\ ,$ (43)
@@
${\cal U}_{\rm exact}(t)\penalty 10000\$$=\penalty 10000\ \left[{\cal U}(\delta t)\right]^{t/\delta t}+{\cal O}(\delta t^{3})\penalty 10000\ .$ (44)
@@
III Initial state preparation
We now turn to the initial state preparation for the entire qumode lattice. We have seen that it is possible to determine the “false vacuum ground state” of a single uncoupled anharmonic qumode using the adiabatic theorem. However the ground state of the QFT is not simply the product of ground states of the uncoupled anharmonic qumode oscillators. As discussed in Ref. [1] the QFT ground state approaches the product of uncoupled qumode ground states for modes in the large mass and low momentum limit, but generally it will involve multi-qumode entanglement and hence long-range correlations. This distinction could be ignored in Ref. [1] because that work was concerned only with scattering processes, so it could accordingly focus on scattering of low momentum wave-packets. Unfortunately for tunnelling processes the distinction is important, because the long-range fluctuations that seed the process are absent from the product of qumode groundstates. In other words, quantum tunnelling could occur on the lattice spontaneously even if it is initialised as the product of qumode ground states, but it could only be a much rarer coherent transition of the whole lattice rather than a localised bubble.
In principle one could initialise the entire lattice using the same adiabatic evolution procedure that we used for a single qumode in Sec. II.2. This would entail starting the system in an MPS which consists of the product of qumodes all in their “false vacuum ground states” determined as described in II.2. Then we would act on the MPS with the following set of $M$ MPOs:
$\widetilde{\mathcal{W}}\penalty 10000\ =\penalty 10000\ \bigotimes_{r=1}^{M}\prod_{n=1}^{N}\left[e^{-i\hat{p}_{n}^{2}\delta t}e^{-i({\mathscr{V}}(\hat{q}_{n})+{\mathscr{V}}_{\rm lift}(\hat{q}_{n}))\delta t}\penalty 10000\ {\widetilde{\mathcal{V}}}^{\,r}_{n}\right]\penalty 10000\ ,$ (47)
@@ III Initial state preparation
$\widetilde{\mathcal{V}}_{n<N}\penalty 10000\ =\penalty 10000\ e^{-i(\hat{q}_{n}^{2}-\hat{q}_{n}\hat{q}_{n+1})\delta s\,\delta t}\penalty 10000\ ,$ (48)
@@
The principle is, of course, exactly the same as for the determination of the “false vacuum ground state” of a single qumode. That is, beginning in an MPS which is a product of qumode oscillator ground states, and adiabatically turning on the space-derivative couplings, which include the hopping terms, should result in the system reaching the “false vacuum ground state” of the QFT. Note, for this purpose that we must add ${\mathscr{V}}_{\rm lift}$ to the potential to ensure that the system does not tunnel during the adiabatic evolution while it is being prepared. Note also that one could combine everything into a single adiabatic procedure, starting in an MPS which is a product of SHO ground states centred at $\varphi_{\rm fv}$, and adiabatically turning on both ${\mathscr{V}}_{1}$ and the space-derivative terms at the same time.
Although appealing from a physics perspective, this kind of procedure has a significant drawback. On a tensor network it would be exceedingly lengthy to implement, because the evolution of the system has to remain adiabatic. Indeed the accurate adiabatic preparation of a single qumode required thousands of Trotter steps, and therefore a currently feasible QFT vacuum preparation by adiabatic evolution would be subject to errors because of loss of adiabaticity.
In this section we present two alternative methods for QFT vacuum preparation. The first is a compromise option: namely we work in the ground state corresponding to a product of uncoupled anharmonic oscillators, and then model the effect of the missing hopping term by sampling. This is an efficient method which is useful for studying phase transitions across the whole system.
The second method is a no-compromise option that requires a separate stage of preparation: namely we find the exact QFT vacuum by imaginary-time TEBD on our qumode lattice. As we shall see this yields the complete QFT “false vacuum ground state”.
III.1 Approximating the ground state by sampling
@@
Let us consider the first method, to approximate the QFT ground state by sampling. For this method one generates an ensemble (labelled by $s$) of initial position and momentum fluctuations, $\{\Delta q_{n}^{(s)},\Delta p_{n}^{(s)}\}$, whose distributions incorporate those quantum fluctuations and correlations of the QFT ground state which are missing when we simply start the system as an MPS which is a product of quantum anharmonic oscillators.
It is straightforward to derive the fluctuations required to simulate the false vacuum of the QFT by sampling a classical distribution. To see how this works, first suppose that there were no quantum fluctuations at all in the system, and that we wished to generate all of the fluctuations by sampling. In that case one would proceed as follows. The fluctuations that we wish to reproduce are encompassed by the position and momentum covariance matrices of the QFT vacuum, $C_{q}$ and $C_{p}$, which are $N\times N$ matrices characterising the correlations between sites:
where $\hat{q}=(\hat{q}_{1},\hat{q}_{2},\ldots,\hat{q}_{N})^{T}$ and similarly for $\hat{p}$. To find a suitable ensemble of $q$ and $p$ values that reproduces these covariance matrices, we Cholesky decompose them as follows:
$C_{q}$ $\penalty 10000\ =\penalty 10000\ L_{q}L_{q}^{T},$
$C_{p}$ $\penalty 10000\ =\penalty 10000\ L_{p}L_{p}^{T},$ (51)
where $L_{q}$ and $L_{p}$ are lower-triangular matrices. It is straightforward to show that an ensemble of displacements and momenta of the following form will reproduce the same covariance matrices (where $s$ simply labels the members of the ensemble):
${q}^{(s)}\penalty 10000\$$=\penalty 10000\ {q}_{\text{mean}}+L_{q}{\eta}^{(s)}_{q}\penalty 10000\$,
${p}^{(s)}\penalty 10000\$$=\penalty 10000\ {p}_{\text{mean}}+L_{p}{\eta}^{(s)}_{p}\penalty 10000\$. (52)
Here $q_{\text{mean}}$ and $p_{\text{mean}}$ are the mean initial displacement and momentum vectors (e.g., $q_{\text{mean}}=\varphi_{\rm fv}$ if starting from the ground state in the false vacuum), and $\eta^{(s)}_{q,p}$ are standard normal random vectors,
${\eta}_{q},{\eta}_{p}\sim\mathcal{N}(\bm{0},\mathbf{I}_{N})\penalty 10000\ ,$ (53)
where ${\eta}^{(s)}_{q},{\eta}^{(s)}_{p}\in\mathbb{R}^{N}$.
The lattice would then be initialised with position and momentum displacements drawn from this ensemble:
  • Positions: $\{q_{j}^{(s)}\}_{j=1}^{N}$
  • Momenta: $\{p_{j}^{(s)}\}_{j=1}^{N}$
and the dynamical simulation can then proceed as normal.
In the current context, as we are interested mainly in missing long-range correlations, it will be sufficient to use the free-field SHO approximation around the false vacuum. This is given by [1]
$\langle{\bf q}|\Psi\rangle\penalty 10000\ \sim\penalty 10000\ \prod_{\alpha}\left({\omega_{\alpha}}^{-\frac{1}{4}}\penalty 10000\ e^{-\omega_{\alpha}|q_{\alpha}|^{2}}\right)\penalty 10000\ ,$ (54)
where the frequencies are given by Eq. (95). With this approximation the computation is considerably simplified. It is simple to show that the covariance matrices are
$C_{q,nm}$ $\penalty 10000\ =\penalty 10000\ \sum_{\alpha}\frac{1}{2\omega_{\alpha}}U^{\dagger}_{n\alpha}U_{\alpha m}$ (55)
$C_{p,nm}$ $\penalty 10000\ =\penalty 10000\ \sum_{\alpha}\frac{\omega_{\alpha}}{2}U^{\dagger}_{n\alpha}U_{\alpha m}\penalty 10000\ ,$ (56)
where $U_{\alpha m}$ is as given in Eq. (92), and one can then directly identify
$L_{q,n\alpha}\penalty 10000\ =\penalty 10000\ \frac{1}{\sqrt{2\omega_{\alpha}}}U^{\dagger}_{n\alpha}\penalty 10000\ ;\penalty 10000\ \penalty 10000\ L_{p,n\alpha}\penalty 10000\ =\penalty 10000\ \sqrt{\frac{\omega_{\alpha}}{2}}\,U^{\dagger}_{n\alpha}\penalty 10000\ .$ (57)
There is only one subtlety in the decomposition, which is the fact that $q_{n}$ and $p_{n}$ are real. Focussing on the position displacements for the moment, one now has
${q}^{(s)}\penalty 10000\ =\penalty 10000\ \varphi_{\rm fv}+\sum_{\alpha=0}^{N-1}\frac{1}{\sqrt{2N\omega_{\alpha}}}e^{-2\pi in\alpha}{\eta}^{(s)}_{q,\alpha}\penalty 10000\$, (58)
${q}_{n}\penalty 10000\ =\penalty 10000\ \varphi_{\rm fv}+\sum_{\alpha=0}^{N-1}\frac{1}{\sqrt{2N\omega_{\alpha}}}\,\cos(2\pi n\alpha/N)\penalty 10000\ {\eta}_{n,\alpha}\penalty 10000\ ,$ (59)
and likewise implement initial momenta of the form
$p_{n}\penalty 10000\ =\penalty 10000\ \sum_{\alpha=0}^{N-1}\sqrt{\frac{\omega_{\alpha}}{2N}}\,\cos(2\pi n\alpha/N)\penalty 10000\ {\eta}^{\prime}_{n,\alpha}\penalty 10000\ .$ (60)
To see how we can identify them, let us first understand a bit better the overall structure of the qumode lattice fluctuations by considering what happens to our result above when there is uniform mode variance, i.e. $\omega_{\alpha}=\omega$. In this case, the fluctuations in both $q$ and $p$ are proportional to $\frac{1}{\sqrt{N}}\sum_{\alpha}\cos(2\pi n\alpha)\eta_{q,\alpha}$. This is essentially the Fourier transform of a white noise vector with independent and identically distributed $\mathcal{N}(0,1)$ entries. It is a straightforward exercise to show that the variance of this sum leads to $\Delta q_{n}\penalty 10000\ =\penalty 10000\ 1/2\omega$ and $\Delta p_{n}\penalty 10000\ =\penalty 10000\ \omega/2$. This variance in the individual qumode quadrature variables is exactly equivalent to the variance of the SHO ground state: in other words, if $\omega_{\alpha}$ were degenerate then the QFT vacuum would be the same as the product of qumode ground states! This is of course entirely in accord with the discussion of Ref. [1] which concluded the two vacua are essentially equivalent if one is concerned with scattering of low-momentum (specifically non-relativistic) modes for which $\omega_{\alpha}\approx m$.
$C_{q,nm}$ $\penalty 10000\ =\penalty 10000\ \sum_{\alpha}\frac{1}{2}\left(\frac{1}{\omega_{\alpha}}-\frac{1}{\omega}\right)\penalty 10000\ U^{\dagger}_{n\alpha}U_{\alpha m}$ (61)
$C_{p,nm}$ $\penalty 10000\ =\penalty 10000\ \sum_{\alpha}\frac{1}{2}\left(\omega_{\alpha}-\omega\right)\penalty 10000\ U^{\dagger}_{n\alpha}U_{\alpha m}\penalty 10000\ .$ (62)